# Cross-Modal Learning for Anomaly Detection in Fused Magnesium Smelting Process: Methodology and Benchmark

### v1

@@ III FmFormer for Anomaly Detection
The proposed FMF Transformer (FmFormer) has two sources of inputs, video with two spatial dimensions and one temporal dimension and three-phase alternating current with one temporal dimension. The FmFormer is composed of a tokenization module, a cross-modal Transformer encoder, and a multi-head decoder. The tokenization module converts 3D video and 1D current into two sets of 1D features. The encoder then explores the correlation between two modalities by using these features. The decoder consists of two functionally different parts, a classification head for class-level anomaly detection and a dense prediction head for spotting anomaly regions, i.e., pixel-level detection. The proposed FmFormer explores the internal features of each modality and the correlation feature across different modalities, achieving state-of-the-art performance in both anomaly detection and anomaly region spotting. The overall architecture of the proposed FmFormer is illustrated in Fig. 2.
@@ III-B1 Internal feature encoding using self-attention
$\phi_{A}(\mathcal{Q}_{f},\mathcal{K}_{f},\mathcal{V}_{f})=\sigma(\frac{\mathcal{Q}_{f}\mathcal{K}_{f}^{T}}{\sqrt{D_{f}}})\mathcal{V}_{f},$ (4)
where $\phi_{A}$ stands for the attention mechanism, $\sigma$ denotes the softmax non-linearity, and $D_{f}$ is the dimension of the input features. To jointly learn different representation subspaces from training instances, the MHSA concatenates multiple self-attentions as follows:
$\begin{split}\mathcal{A}_{f}&=\{(A_{f}^{(1)})^{T},\dots,(A_{f}^{(L_{s})})^{T}\}^{T}\mathcal{W}_{f}^{s},\\ A_{f}^{(l_{s})}&=\phi_{A}(\mathcal{Q}^{(l_{s})}_{f},\mathcal{K}^{(l_{s})}_{f},\mathcal{V}^{(l_{s})}_{f}),l_{s}\in\{1,\cdots,L_{s}\}\\ \mathcal{Q}_{f}&=\{(\mathcal{Q}^{(1)}_{f})^{T},\dots,(\mathcal{Q}^{(L_{s})}_{f})^{T}\}^{T},\\ \mathcal{K}_{f}&=\{(\mathcal{K}^{(1)}_{f})^{T},\dots,(\mathcal{K}^{(L_{s})}_{f})^{T}\}^{T},\\ \mathcal{V}_{f}&=\{(\mathcal{V}^{(1)}_{f})^{T},\dots,(\mathcal{V}^{(L_{s})}_{f})^{T}\}^{T},\end{split}$ (5)
where $A_{f}^{(l_{s})}$ denotes one of the self-attention head, $L_{s}$ is the number of self-attention in the MHSA, $\mathcal{W}_{f}^{s}\in\mathbb{R}^{D\times D}$ stands for a weight matrix to linearly map the concatenated self-attention features into dimension $D$, and $\mathcal{A}_{f}\in\mathbb{R}^{N_{f}\times D}$ is the resulting MHSA feature. In (5), $\mathcal{Q}_{f}$, $\mathcal{K}_{f}$ and $\mathcal{V}_{f}$ are evenly divided into $\mathcal{Q}^{(l_{s})}_{f},\mathcal{K}^{(l_{s})}_{f},\mathcal{V}^{(l_{s})}_{f}\in\mathbb{R}^{N_{f}\times\frac{D}{L_{s}}}$ for the computation of each self-attention head. The final tokens processed by the MHSA is:
$\begin{split}z^{\prime}_{f}&=\mathcal{A}_{f}+z_{f},\\ z_{f}&\leftarrow\phi_{MLP}(z^{\prime}_{f})+z^{\prime}_{f},\end{split}$ (6)
@@ III-B2 Correlation feature encoding using bidirectional cross-attention
To facilitate feature exploration across different modalities, we utilize the cross-attention between visual and current modalities in a bidirectional manner. The implemented cross-attention has the following characteristics: i) Each token in one modality can interact with all the tokens in the other modality; ii) The resulting attention map is directional sensitive; iii) The dimension of each modality remain unchanged.
@@
$\begin{split}z^{\prime}_{c\rightarrow v}&=\mathcal{A}_{c\rightarrow v}+z_{v},\\ z_{c\rightarrow v}&\leftarrow\phi_{MLP}(z^{\prime}_{c\rightarrow v})+z^{\prime}_{c\rightarrow v},\\ z^{\prime}_{v\rightarrow c}&=\mathcal{A}_{v\rightarrow c}+z_{c},\\ z_{v\rightarrow c}&\leftarrow\phi_{MLP}(z^{\prime}_{v\rightarrow c})+z^{\prime}_{v\rightarrow c},\end{split}$ (8)
@@
It should be noted that the cross-attention map between the two modalities in (7) is directional sensitive, i.e., $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{v}(\mathcal{K}^{(l_{c})}_{c})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{v}\times N_{c}}$ and $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{c}(\mathcal{K}^{(l_{c})}_{v})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{c}\times N_{v}}$. Besides, it encodes the correlation of each token in one modality with all the other tokens in the other modality. Therefore, the bidirectional cross-attention enables the proposed FmFormer to have a global perception across modalities.
@@ III-C1 Dense prediction head
where $\hat{y}^{pix}\in\mathbb{R}^{H\times W\times K}$ denotes the pixel-level prediction result and $K=2$ indicates the possibility of normal and abnormal.
@@ III-C2 Classification head
The classification head transforms the class tokens into an anomaly prediction, i.e., a classification vector. First, we map the class tokens (marked in dark green and dark violet in Fig. 2(c)) processed by the Transformer encoder, $z_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times D}$ and $z_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times D}$, to vectors $\hat{y}_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times K}$ and $\hat{y}_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times K}$ using two small MLPs. Then the detection result is obtained via a simple linear fusion mechanism formulated as follows:
@@ IV-A Fused Magnesium Smelting Process Benchmark
The benchmark contains cross-modal data from fused magnesium smelting processes for a total of 3 production batches, in which the video data is captured by industrial cameras, and the current data is sampled by PLC control systems, as shown on the right of Fig. 1(a). Signal generators are employed to trigger industrial cameras and PLC control systems to capture video frames and currents, thus ensuring synchronous acquisition. The raw data contains a total of more than 1,000 hours of synchronized videos and current sequences. However, directly using the raw dataset as the benchmark causes class imbalance, as abnormal conditions are infrequent. Therefore, we carefully selected over $2.2\times 10^{6}$ samples (about 25 hours, 25 samples per second, and spatial resolution of $1440\times 2560$ for each raw video frame) through the entire raw dataset to keep the ratio of normal to abnormal samples close to $1:1$. The benchmark comprises two types of datasets: a pixel-level annotated dataset with approximately $2.7\times 10^{5}$ samples and a class-level annotated dataset with around $2.0\times 10^{6}$ samples. More detailed statistics for the benchmark are listed in Table I. The fused magnesium smelting process benchmark will be available at https://github.com/GaochangWu/FMF-Benchmark.
@@
Figure 3: Efficient data annotation based on the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. (a) The initial annotation is achieved by the following steps: i) determining the onset, apex and offset times of anomalies according to the features summarized by domain experts; ii) labelling the sparse set of frames (keyframes) with boxes; iii) propagating boxes to other frames using their adjacent keyframes. (b) The annotation refinement is implemented by applying a weighted median filter to each label, providing a more accurate representation of the anomaly region.
The annotation of the large volume cross-modal data requires a mass of human labour. To accelerate the data annotation, we take advantage of the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. More specifically, despite the data can be disturbed by noise or visual occlusion, the actual anomaly remains constant for a certain period of time. Based on this observation, we first determine the onset (starting), apex (highest intensity) and offset (ending) times of anomalies according to the visual and current features summarized by domain experts, and coarsely annotate a sparse set of frames using boxes, each of which is called a keyframe. A more detailed description of the normal/abnormal current features can be referred in [3, 2, 1]. The coarse labels are then propagated to other frames between each pair of adjacent keyframes, providing every frame a coarse label, as shown in Fig. 3(a). To obtain more accurate labels for the pixel-level anomaly detection, we adopt a weighted median filter [57] to refine each label guided by the corresponding frame, as shown in Fig. 3(b). With the help of the efficient data annotation, the presented benchmark provides approximately $2.7\times 10^{5}$ cross-modal samples with pixel-level labels. Fig. 4 demonstrates some of the examples of video frames (superimposed with labels) and current curves in the presented cross-modal benchmark.
@@ IV-B Training Recipe
where $\mathcal{W}$ and $e$ are the learnable weights and learnable positional embeddings in the proposed FmFormer, $\langle x_{v},x_{c},y\rangle$ is the training set of input video $x_{v}$, input current $x_{c}$, and label $y$ triples, $\mathcal{L}^{pix}$ is the dense prediction loss term, $\mathcal{L}^{cls}$ is the classification loss term, and $\alpha$ is a hyperparameter to control the balance of the two terms. Specifically, we use the pixel-level cross-entropy loss for the dense prediction:
@@
where $(h,w,k)$ indicates an element index in the prediction result $\hat{y}^{pix}$ or the pixel-label $y$. For the classification loss, we first aggregate the pixel-level labels $y$ to the class-level label by determining whether there is anomaly, i.e., the type is classified as abnormal if the label $y$ contains any anomaly region; otherwise, it is categorized as normal. Then we use class-wise cross-entropy loss for the classification:
$\begin{split}\mathcal{L}^{cls}(\hat{y}^{cls},y)&=-\sum_{k\in K}\mathcal{G}(y)_{(k)}\log\hat{y}^{cls}_{(k)},\\ \mathcal{G}(y)&=\begin{cases}[0,1],&\text{if }\sum_{h,w}y_{(h,w,2)}>\tau\\ [1,0],&\text{otherwise},\end{cases}\end{split}$ (13)
where $\mathcal{G}(\cdot)$ denotes the aggregation operation that converts pixel-level labels into class-level labels as described above, and $\tau$ is a threshold which we empirically set to $0.5$.
@@ IV-C1 Architecture details
We use a kernel of size $2\times 8\times 8$ (time $t_{v}$, height $h_{v}$, and width $h_{w}$) and a kernel of dilation rate of 2 with the same kernel size for the video tokenization. Throughout the proposed FmFormer, each token has a dimension of $D=96$. In the cross-modal Transformer encoder, the numbers of heads in the MHSA and MHCA are set to $L_{s}=3$ and $L_{c}=3$, respectively. We alternately cascade $L_{e}=6$ layers of MHSA and MHCA as shown in Fig. 2(b). In the dense prediction head, the $1\times 1$ convolution layers in (9) have channel number $D=96$, and each transposed convolution layers in (10) and (11) has a $2\times 2$ kernel size with stride $2\times 2$. For simplicity, each convolution layer in (10) and (11) is practically a sequential connection of a $3\times 3$ convolution, a batch normalization, and a ReLU non-linearity. Due to the spatial reconstruction, we set the channel numbers of the transposed convolution layers in (11) as $[128,64,32,\cdots]$. The MLPs in the MHSA, MHCA (in (6) and (8)) and the classification head are sequential connections of a layer normalization, a linear layer, a GELU non-linearity, and another linear layer.
@@ IV-C2 Training details
For the joint training of the dense prediction head and the classification head, we use the pixel-level annotated dataset as described in Section IV-A. Note that pixel-level labels can also be converted into class-level labels. The dataset is divided into a training set with about $2.15\times 10^{5}$ examples and a test dataset with $0.55\times 10^{5}$ examples. In both the training and test datasets, the ratio of normal examples to abnormal examples remained close to $1:1$. We crop out areas of interest (i.e., furnaces) from the raw videos for the training and testing. The proposed FmFormer is implemented by using the Pytorch framework [58]. The AdamW solver [59] is applied as the optimization method, in which the batch size is set to 64. The step learning rate decay scheme is adopted with an initial learning rate of $5\times 10^{-4}$, which then decays to $5\times 10^{-5}$ after 20 epochs. The network converges after 30 epochs of training, which takes about 5 hours on an NVIDIA TESLA V100.

### v2

@@ III FmFormer for Anomaly Detection
The proposed FMF Transformer (FmFormer) has two sources of inputs, video with two spatial dimensions and one temporal dimension and three-phase alternating current with one temporal dimension. The FmFormer is composed of a tokenization module, a cross-modal Transformer encoder, and a multi-head decoder. The tokenization module converts 3D video and 1D current into two sets of 1D features. The encoder then explores the correlation between two modalities by using these features. The decoder consists of two functionally different parts, a classification head for class-level anomaly detection and a dense prediction head for spotting anomaly regions, i.e., pixel-level detection. The overall architecture of the proposed FmFormer is illustrated in Fig. 2.
@@ III-B1 Internal feature encoding using self-attention
$A_{f}^{(l_{s})}=\phi_{A}(\mathcal{Q}_{f}^{(l_{s})},\mathcal{K}_{f}^{(l_{s})},\mathcal{V}_{f}^{(l_{s})})=\sigma(\frac{\mathcal{Q}_{f}^{(l_{s})}(\mathcal{K}_{f}^{(l_{s})})^{T}}{\sqrt{D/L_{s}}})\mathcal{V}_{f}^{(l_{s})},$ (4)
where $A_{f}^{(l_{s})}$ denotes one of the attention heads, $l_{s}\in\{1,\cdots,L_{s}\}$ with $L_{s}$ being the number of attention in the MHSA, $\phi_{A}$ stands for the attention mechanism, and $\sigma$ denotes the softmax non-linearity. Here, $\mathcal{Q}_{f}$, $\mathcal{K}_{f}$ and $\mathcal{V}_{f}$ are evenly divided into $\mathcal{Q}^{(l_{s})}_{f},\mathcal{K}^{(l_{s})}_{f},\mathcal{V}^{(l_{s})}_{f}\in\mathbb{R}^{N_{f}\times\frac{D}{L_{s}}}$ for the computation of each self-attention head. To jointly learn different representation subspaces from training instances, the MHSA concatenates multiple self-attention results as follows:
$\mathcal{A}_{f}=\{(A_{f}^{(1)})^{T},\dots,(A_{f}^{(L_{s})})^{T}\}^{T}\mathcal{W}_{f}^{s}+z_{f},$ (5)
where $\mathcal{W}_{f}^{s}\in\mathbb{R}^{D\times D}$ stands for a weight matrix to linearly map the concatenated self-attention features into dimension $D$, and $\mathcal{A}_{f}\in\mathbb{R}^{N_{f}\times D}$ is the resulting MHSA feature. The final tokens processed by the MHSA is:
$z_{f}\leftarrow\phi_{MLP}(\mathcal{A}_{f})+\mathcal{A}_{f},$ (6)
@@ III-B2 Correlation feature encoding using bidirectional cross-attention
To facilitate feature exploration across different modalities, we utilize the cross-attention between visual and current modalities in a bidirectional manner. The implemented cross-attention has the following characteristics: i) Each token in one modality can interact with all the tokens in the other modality; ii) The resulting attention map has two directions, current-to-visual and visual-to-current, and is directional sensitive; iii) The dimension of each modality remain unchanged.
@@
$\begin{split}z_{c\rightarrow v}&\leftarrow\phi_{MLP}(\mathcal{A}_{c\rightarrow v})+\mathcal{A}_{c\rightarrow v},\\ z_{v\rightarrow c}&\leftarrow\phi_{MLP}(\mathcal{A}_{v\rightarrow c})+\mathcal{A}_{v\rightarrow c},\end{split}$ (8)
@@
It should be noted that the MHCA in (7) is directional sensitive, i.e., the attention maps $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{v}(\mathcal{K}^{(l_{c})}_{c})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{v}\times N_{c}}$ and $\sigma(\frac{\mathcal{Q}^{(l_{c})}_{c}(\mathcal{K}^{(l_{c})}_{v})^{T}}{\sqrt{D/L_{c}}})\in\mathbb{R}^{N_{c}\times N_{v}}$ hold different meanings based on the direction of the information flow between the visual and current modalities. Specifically, in current-to-visual directional MHCA, the visual features query the current features, producing an attention map that encodes the correlation of each visual token according to its attention from all current tokens, and vice versa. Therefore, the bidirectional cross-attention enables the proposed FmFormer to achieve global perception across modalities. This also ensures that the model captures more comprehensive features, enabling accurate and robust anomaly detection in the fused magnesium smelting process, particularly when one of the modality are less reliable.
@@ III-C1 Dense prediction head
where $\hat{y}^{pix}\in\mathbb{R}^{H\times W\times K}$ denotes the pixel-level prediction result and $K=2$ indicates the binary classification task (normal and abnormal).
@@ III-C2 Classification head
The classification head transforms the class tokens into an anomaly prediction, i.e., a classification vector. First, we map the class tokens (marked in dark green and dark violet in Fig. 2(c)) processed by the Transformer encoder, $z_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times D}$ and $z_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times D}$, to vectors $\hat{y}_{c\rightarrow v}^{cls}\in\mathbb{R}^{1\times K}$ and $\hat{y}_{v\rightarrow c}^{cls}\in\mathbb{R}^{1\times K}$ using two small MLPs. Then the detection result is obtained via a simple Linear Fusion (LF) mechanism formulated as follows:
@@ IV-A Fused Magnesium Smelting Process Benchmark
The benchmark contains cross-modal data from fused magnesium smelting processes for a total of 3 production batches, in which the video data is captured by industrial cameras, and the current data is sampled by PLC control systems, as shown on the right of Fig. 1(a). Rectangular wave signal generators are employed to trigger industrial cameras and PLC control systems to capture video frames and currents, enabling the synchronous acquisition of video frames and current measurements. The raw data contains a total of more than 1,000 hours of synchronized videos and current sequences. However, directly using the raw dataset as the benchmark causes class imbalance, as abnormal conditions are infrequent. Therefore, we carefully selected over $2.2\times 10^{6}$ samples (about 25 hours, 25 samples per second, and spatial resolution of $1440\times 2560$ for each raw video frame) through the entire raw dataset to keep the ratio of normal to abnormal samples close to $1:1$. The benchmark comprises two types of datasets: a pixel-level annotated dataset with approximately $2.7\times 10^{5}$ samples and a class-level annotated dataset with around $2.0\times 10^{6}$ samples. More detailed statistics for the benchmark are listed in Table I. The fused magnesium smelting process benchmark will be available at https://github.com/GaochangWu/FMF-Benchmark.
@@
Figure 3: Efficient data annotation based on the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. (a) The initial annotation is achieved by the following steps: i) determining the onset, apex and offset times of anomalies; ii) labelling the sparse set of frames (keyframes) with bounding boxes; iii) propagating boxes to other frames. (b) The annotation refinement is implemented to provide a more accurate representation of the anomaly region.
IV-A1 Data annotation
The annotation of the large volume cross-modal data requires a mass of manpower. To accelerate the data annotation, we take advantage of the piecewise temporal consistency of abnormal conditions in fused magnesium smelting processes. Specifically, despite the data can be disturbed by noise or visual occlusion, the actual anomaly remains constant for a certain period of time. Based on this observation, we first determine the onset (starting), apex (highest intensity) and offset (ending) times of anomalies according to the visual and current features summarized by domain experts, and coarsely annotate a sparse set of frames using bounding boxes, each of which is called a keyframe. A more detailed description of the normal/abnormal current features can be referred in [3, 2, 1]. Next, a label propagation process is employed to assign a coarse label to every frame, as shown in Fig. 3(a). It linearly projects the bounding boxes to each frame between onset keyframe to apex keyframe, and then to each frame between apex keyframe and offset keyframe. To obtain more accurate labels that better fit the shape of the anomalies for dense prediction, we finally implement an annotation refinement process using a weighted median filter [55], as shown in Fig. 3(b). It refines each label by utilizing the corresponding video frame as guidance. With the help of the efficient data annotation, the presented benchmark provides approximately $2.7\times 10^{5}$ cross-modal samples with pixel-level labels.
IV-A2 Analysis of data modality
Fig. 4 demonstrates examples of video frames (superimposed with labels) and current curves from the presented cross-modal benchmark. The first row illustrates four normal cases, while the last two rows show eight abnormal cases. From the first and last rows, it is evident that abnormal patterns in current signals can be easily obscured by data noise, complicating anomaly detection when relying solely on the current modality. In comparison, despite visual information showing salient features in most cases, it can be affected by occlusion from heavy water mist, as shown in the second and fourth cases of the last row. Additionally, weak feature anomalies can be obscured by intense flame light, as demonstrated by the second and third cases in the second row. Our benchmark provides both visual and current information to address these challenges and enhance the robustness of anomaly detection in fused magnesium smelting processes.
@@ IV-B Training Recipe
where $\mathcal{W}$ and $e$ are the learnable weights and learnable positional embeddings in the proposed FmFormer, $\langle x_{v},x_{c},y\rangle$ is the training set of input video $x_{v}\in\mathbb{R}^{T_{v}\times H\times W\times 3}$, input current $x_{c}\in\mathbb{R}^{T_{c}\times 3}$, and label $y\in\mathbb{R}^{H\times W\times K}$ triples, $\mathcal{L}^{pix}$ is the dense prediction loss term, $\mathcal{L}^{cls}$ is the classification loss term, and $\alpha$ is a hyperparameter to control the balance of the two terms. Specifically, we use the pixel-level cross-entropy loss for the dense prediction:
@@
where $(h,w,k)$ indicates an element index in the prediction result $\hat{y}^{pix}$ or the pixel-label $y$. For the classification loss, we first employ an aggregation operation $\mathcal{G}$ to convert the pixel-level labels $y$ to the class-level label, simply by determining the presence of anomalies. Then we use class-wise cross-entropy loss for the classification:
$\mathcal{L}^{cls}(\hat{y}^{cls},y)$ $=-\sum_{k\in K}\mathcal{G}(y;\tau)\log\hat{y}^{cls}_{(k)},$ (13)
where $\mathcal{G}(\cdot;\tau)=[0,1]$ when the number of anomaly pixels in the label $y$ exceeds $\tau$, and $\mathcal{G}(\cdot;\tau)=[1,0]$ otherwise. $\tau$ is a threshold which we empirically set to $0.5$.
@@ IV-C1 Architecture details
Following typical Transformer backbones [14, 16, 28], we implement four FmFormer models with varying capacities (Tiny, Small, Base, Large) with detailed configurations in Table II. In the table, “Token dim.” refers to the dimension of tokens, “MLP dim.” represents the dimension of the MLP in (6), “MHSA heads #” indicates the number of attention heads ($L_{s}$ and $L_{c}$), and “Encoder layers #” denotes the number of Transformer encoder layers. We maintain the same number of layers as in ViT [16] for each capacity configuration (noting that each layer contain one MHSA and one MHCA in our model), while adopting smaller token and MLP dimensions. This makes the model relatively lightweight compared to commonly used large-scale models [16, 28], ensuring computationally efficient with satisfactory performance for industrial deployment, as analyzed in Sect. V-B2.
In the tokenization, we use a kernel of size $2\times 8\times 8$ (time $t_{v}$, height $h_{v}$, and width $h_{w}$) for the video tokenization. This setting is inspired by ViViT [28], with a smaller spatial size to accommodate our dilated tokenization, which incorporates a dilation rate of 2. In the dense prediction head, the $1\times 1$ convolution layers in (9) have the same number of channels as the token dimension $D$, while each transposed convolution layers in (10) and (11) has a $2\times 2$ kernel size with stride $2\times 2$, progressively enhancing the spatial resolution of the feature maps. Following the implementation from [52], each convolution layer in (10) and (11) is practically a sequential connection of a $3\times 3$ convolution, a batch normalization, and a ReLU non-linearity. Due to the spatial reconstruction, we empirically set the channel numbers of the transposed convolution layers in (11) as $[128,64,32,\cdots]$. To align with existing vision Transformers [16, 28], the MLPs in the MHSA, MHCA (in (6) and (8)) and the classification head are sequential connections of a layer normalization, a linear layer, a GELU non-linearity, and another linear layer.
@@ IV-C2 Training details
For the joint training of the dense prediction head and the classification head, we use the pixel-level annotated dataset as described in Section IV-A. Note that pixel-level labels can also be converted into class-level labels. The dataset is divided into a training set with about $2.15\times 10^{5}$ examples and a test dataset with $0.55\times 10^{5}$ examples. In both the training and test datasets, the ratio of normal examples to abnormal examples remained close to $1:1$. We crop out areas of interest (i.e., furnaces) from the raw videos for the training and testing. The proposed FmFormer is implemented by using the Pytorch framework [56]. The AdamW solver [57] is applied as the optimization method, in which the batch size is set to 64. The step learning rate decay scheme is adopted with an initial learning rate of $5\times 10^{-4}$, which then decays to $5\times 10^{-5}$ after 20 epochs. The network (FmFormer-B) converges after 30 epochs of training, which takes about 5 hours on an NVIDIA TESLA V100.
TABLE II: Configuration for each FmFormer model.
Name Token dim. MLP dim. MHSA heads # Encoder layers # Param. #
Tiny 36 144 3 6 0.28M
Small 48 192 3 6 0.48M
Base 96 384 3 6 1.86M
Large 96 384 3 12 3.64M